# Stronger Baselines for Retrieval-Augmented Generation with Long-Context Language Models

### v1

@@ Abstract
With the rise of long-context language models (LMs) capable of processing tens of thousands of tokens in a single pass, do multi-stage retrieval-augmented generation (RAG) pipelines still offer measurable benefits over simpler, single-stage approaches? To assess this question, we conduct a controlled evaluation for QA tasks under systematically scaled token budgets, comparing two recent multi-stage pipelines, ReadAgent and RAPTOR, against three baselines, including DOS RAG (Document’s Original Structure RAG), a simple retrieve-then-read method that preserves original passage order. Despite its straightforward design, DOS RAG consistently matches or outperforms more intricate methods on multiple long-context QA benchmarks. We recommend establishing DOS RAG as a simple yet strong baseline for future RAG evaluations, pairing it with emerging embedding and language models to assess trade-offs between complexity and effectiveness as model capabilities evolve.111We release our code at https://github.com/alex-laitenberger/stronger-baselines-rag.
@@
Alex Laitenberger and Christopher D. Manning and Nelson F. Liu Stanford University, USA alaiten@stanford.edu
@@ 1 Introduction
Recent advances in long-context language models (LMs), such as GPT-4o, have expanded the token processing capabilities of LMs, enabling them to handle tens of thousands of tokens in a single pass. This raises a pivotal question: Are complex, multi-stage retrieval-augmented generation (RAG) pipelines still necessary when simpler, single-stage methods can now leverage these extended contexts effectively?
Retrieval-augmented generation (RAG) systems traditionally combine a retriever, which selects passages from a large corpus relevant to a given query, and a reader, typically a language model, to generate a final answer. (Lewis et al., 2020). Prior work has proposed a variety of retrieval strategies to circumvent the limited long-context reasoning ability of earlier reader LMs. For example, abstractive preprocessing, iterative chunk summarization, and agent-based retrieval loops have been used to compress or reason over texts that might otherwise exceed the input limits of early LMs (Chen et al., 2023; Sarthi et al., 2024; Lee et al., 2024; Sun et al., 2024, inter alia). While effective, these pipelines often introduce significant complexity and computational overhead.
In contrast, modern long-context LMs can now process substantial input sizes directly, suggesting that simpler retrieve-then-read strategies might suffice in certain settings. To assess the relevance of multi-stage pipelines under these new conditions, we conduct a controlled evaluation in which we systematically increase token budgets, analyzing how effectively each approach leverages extended contexts using a modern long-context LM (GPT-4o) as the downstream reader. We compare two recent multi-stage pipelines (ReadAgent and RAPTOR; Lee et al., 2024; Sarthi et al., 2024) against three baselines, including DOS RAG (Document’s Original Structure RAG).
DOS RAG maintains a simple retrieve-then-read strategy and presents retrieved passages in their original document order. Despite its simplicity, our findings across three QA benchmarks (Bench, QuALITY, NarrativeQA) indicate that DOS RAG can match or even outperform more complex multi-stage pipelines on all evaluated retrieval token budgets (Figure 1, Figure 2, Figure 3).
This work advocates for establishing DOS RAG as a simple yet strong baseline for RAG evaluations, pairing it with emerging embedding and language models to benchmark trade-offs between complexity and effectiveness as model capabilities continue to evolve.
@@ 2 Experimental Setup
We compare the performance of two recent multi-stage RAG pipelines (ReadAgent and RAPTOR) against three baselines (Vanilla RAG, the full-document baseline, and DOS RAG) on three long-context question-answering tasks (Bench, QuALITY and NarrativeQA). See Appendix A for further details about experimental setup, implementation, and used prompts.
@@ 2.1 Benchmarks
The NarrativeQA benchmark is a long-document question-answering dataset in English that challenges models to answer questions about stories by reading entire books or movie scripts (Kočiský et al., 2018). We evaluate on the test set, which contains 355 stories (average length of 57K tokens) and 10,557 questions. Each story’s questions are constructed such that high performance requires understanding the underlying narrative, versus relying on shallow pattern matching.
@@ 2.2 Multi-Stage RAG Pipelines
RAPTOR handles long documents by recursively organizing the text into a tree of hierarchical summaries (Sarthi et al., 2024). Concretely, it partitions the text into sentence-level chunks, clusters related chunks, and uses a language model to summarize each cluster. This process repeats, generating higher-level summaries until a final set of root nodes represents the entire document. At inference time, RAPTOR retrieves from different levels of the summary tree, balancing broad coverage against local detail.
@@ 2.3 Baselines
In our implementation the corpus is first split into chunks capped at 100 tokens, while preserving sentence boundaries where possible. We use neural retrieval with a sentence embedding model (Snowflake Arctic-embed-m 1.5 by Merrick, 2024) to encode both the query and the resulting text chunks into a shared embedding space. At inference time, chunks are ranked by cosine similarity to the query embedding, with the top-ranked chunks retrieved until a fixed input token budget (e.g., 10K tokens) is reached. The selected chunks, ordered by decreasing similarity, are then concatenated with the query to construct the input to the language model.
@@
The full-document baseline prompts the model with all available text, preserving passage continuity and reducing the risk of missing key information through retrieval errors. We evaluate it only on the QuALITY benchmark, where document lengths fit within the model’s context window.
@@
DOS RAG follows the same retrieval and embedding process as Vanilla RAG, but with one key difference: retrieved chunks are reordered to match their original document order, not sorted by similarity score. This reordering, achieved by tracking chunk positions, preserves passage continuity like the full-document baseline while still filtering irrelevant content like Vanilla RAG.
@@ 3 Results
Figure 2 shows performance on the QuALITY benchmark, again with GPT-4o as the reader model. In this setting, we see that all approaches see a steady rise in accuracy as the retrieval budget grows. In particular, full-document baseline with GPT-4o achieves 91.2%, outperforming the best retrieval-augmented systems. Among the retrieval-augmented methods, DOS RAG again achieves the highest performance for token budgets of up to 8K.
@@ 4 Analysis
Why is DOS RAG outperforming other examined methods?
We identify four key strategies that underlie DOS RAG’s performance and align with empirical findings from our evaluation:
1. Retrieving from original passages rather than generated summaries, thereby preserving source information, as in Vanilla RAG and the full-document baseline: We implemented Vanilla RAG as an exact ablation of RAPTOR that excludes generated summaries from the retrieval process. Vanilla RAG consistently outperforms RAPTOR across datasets and retrieval sizes, reinforcing our hypothesis that retrieving directly from original passages results in more robust QA, particularly as long-context LMs reduce the need for intermediate abstraction.
2. Prioritizing retrieval recall over precision while staying within the LM’s effective context size: DOS RAG’s performance increases consistently as the retrieval budget expands up to 30K tokens, after which it plateaus and declines, aligning with prior findings that LMs’ effective context length remains limited (Liu et al., 2024). For shorter documents (6k–8k tokens), the full-document baseline outperforms all methods, indicating that maximizing recall, by including critical information anywhere in the input, can be more effective than precision filtering. However, beyond the effective context window, eliminating irrelevant chunks remains essential to maintain performance.
3. Reordering retrieved passages to maintain narrative and argument continuity: Vanilla RAG serves as an exact ablation of DOS RAG, excluding the reordering step. Across all benchmarks and retrieval budgets, DOS RAG consistently outperforms Vanilla RAG, underscoring the benefits of preserving passage order. Performance gain is especially high when the retrieval budget is expanded to tens of thousands of tokens. Retrieving more chunks brings us closer to the original document, but without order, the input becomes a disjointed, shuffled version.
4. Simple vs. complex: Multi-stage, agentic approaches like ReadAgent decompose QA into multiple LM calls, increasing token usage and latency. However, our evaluation shows that this added complexity does not necessarily improve performance. ReadAgent underperforms compared to DOS RAG at lower token budgets, highlighting the effectiveness of simpler RAG pipelines that use strong embedding models and long-context LMs.
@@ 6 Conclusion
This work examined whether complex multi-stage retrieval pipelines still justify their added complexity given the emergence of long-context LMs capable of processing tens of thousands of tokens. Our controlled evaluation under systematically scaled token budgets shows that simpler methods like DOS RAG can effectively match or even outperform multi-stage pipelines such as ReadAgent and RAPTOR in QA tasks, without intermediate summarization or agentic processing. Based on these findings, we recommend establishing DOS RAG as a simple yet strong baseline for future RAG evaluations, pairing it with emerging embedding and language models to systematically assess trade-offs between complexity and effectiveness as model capabilities evolve.
@@
Although our results indicate that simpler retrieval-then-read approaches can match or outperform more intricate multi-stage RAG pipelines when paired with long-context language models, our study has several limitations that qualify the generality of these findings.
Our experiments focus on multiple-choice and short-answer reading comprehension tasks on single long input documents. While these settings provide useful testbeds for long-context reasoning, it is unclear whether the trends hold for more diverse tasks such as open-ended generation, tasks that require reasoning over multiple documents, or complex reasoning that requires specialized domain knowledge (e.g., in scientific or legal domains). Future work should investigate whether the benefits of simply preserving document continuity extend to these settings and whether specialized retrieval or summarization steps prove more valuable.
In addition, cost and latency considerations factor heavily into whether large-context prompting offers a practical alternative to complex multi-stage RAG pipelines in practice. While our comparisons accounted for comparable input token budgets during inference, they did not comprehensively assess efficiency across the full processing pipeline, including the costs of embedding and preprocessing documents. We estimate that more complex preprocessing translates into higher cost, but further work is needed to quantify these trade-offs systematically, particularly for high-throughput or resource-limited settings.
@@ A.7 RAPTOR
We implement RAPTOR using the official repository. To match our other systems, we use the NLTK library for sentence segmentation and the Snowflake Arctic-embed-m 1.5 embedding model (Merrick, 2024) to embed and cluster text chunks. In all experiments, we use GPT-4o-mini to build the tree of hierarchical summaries to reduce API costs (though note that we experiment with both GPT-4o-mini and GPT-4o as the downstream reader).
@@ Appendix C Full Results
Table 5 presents the results for the NarrativeQA test set across various systems and baselines, using GPT-4o-mini as the reader. Due to issues with the original NarrativeQA download script, three out of 355 stories from the test set were inaccessible, as their document files were empty. Consequently, our results are reported for 352 documents and 10,391 questions for all methods.
@@
| Avg Spent / Budget | F1 | Bleu-1 | Bleu-4 | Rogue-L | Meteor | |

### v2

@@ Abstract
With the rise of long-context language models (LMs) capable of processing tens of thousands of tokens in a single context window, do multi-stage retrieval-augmented generation (RAG) pipelines still offer measurable benefits over simpler, single-stage approaches? To assess this question, we conduct a controlled evaluation for QA tasks under systematically scaled token budgets, comparing two recent multi-stage pipelines, ReadAgent and RAPTOR, against three baselines, including DOS RAG (Document’s Original Structure RAG), a simple retrieve-then-read method that preserves original passage order. Despite its straightforward design, DOS RAG consistently matches or outperforms more intricate methods on multiple long-context QA benchmarks. We trace this strength to a combination of maintaining source fidelity and document structure, prioritizing recall within effective context windows, and favoring simplicity over added pipeline complexity. We recommend establishing DOS RAG as a simple yet strong baseline for future RAG evaluations, paired with state-of-the-art embedding and language models, and benchmarked under matched token budgets, to ensure that added pipeline complexity is justified by clear performance gains as models continue to improve.111We release our code at https://github.com/alex-laitenberger/stronger-baselines-rag.
@@
Alex Laitenberger and Christopher D. Manning and Nelson F. Liu Stanford University, USA alex.laitenberger@gmail.com / alaiten@stanford.edu, {manning, nfliu}@cs.stanford.edu
@@ 1 Introduction
Recent advances in long-context language models (LMs) have expanded their token processing capabilities, enabling them to handle tens of thousands of tokens in a single context window. This raises a pivotal question: Are complex, multi-stage retrieval-augmented generation (RAG) pipelines still necessary when simpler, single-stage methods can now leverage these extended contexts effectively?
RAG systems traditionally combine a retriever, which selects passages from a large corpus relevant to a given query, and a reader, typically an LM, to generate a final answer (Lewis et al., 2020). de Prior work has proposed a variety of complex, multi-stage retrieval strategies to circumvent the limited long-context reasoning ability of earlier reader LMs. For example, abstractive preprocessing, iterative passage summarization, and agent-based retrieval loops have been used to compress or reason over documents that might otherwise exceed the input limits of early LMs (Chen et al., 2023; Sarthi et al., 2024; Lee et al., 2024; Sun et al., 2024, inter alia). While effective, these pipelines often introduce significant complexity and computational overhead.
In contrast, modern long-context LMs can now directly process substantial amounts of text, suggesting that simpler retrieve-then-read strategies might suffice in certain settings. To compare multi-stage pipelines vs. simpler retrieve-then-read strategies, we conduct a controlled evaluation in which we systematically increase token budgets, analyzing how effectively each approach leverages extended contexts using a representative modern long-context LM (GPT-4o) as the downstream reader (see §2). We compare two recent multi-stage pipelines (ReadAgent and RAPTOR; Lee et al., 2024; Sarthi et al., 2024) against three baselines, including DOS RAG (Document’s Original Structure RAG). DOS RAG maintains a simple retrieve-then-read strategy and presents retrieved passages in their original document order. Despite its simplicity, our findings across three QA benchmarks (Bench, QuALITY, NarrativeQA) indicate that DOS RAG can match or even outperform more complex multi-stage pipelines on all evaluated retrieval token budgets (see §3). Our analysis suggests that DOS RAG’s strength lies in preserving original passages and document structure, prioritizing recall within effective context windows, and maintaining simplicity over added pipeline complexity (see §4).
This work advocates for establishing DOS RAG as a simple yet strong baseline for RAG evaluations, paired with state-of-the-art embedding and language models and benchmarked under matched token budgets, so that added pipeline complexity is justified only when it delivers clear performance gains as model capabilities continue to evolve.
@@ 2 Experimental Setup
We compare the performance of two recent multi-stage RAG pipelines (ReadAgent and RAPTOR) against three baselines (Vanilla RAG, the full-document baseline, and DOS RAG) on three long-context question-answering tasks (Bench, QuALITY and NarrativeQA). See Figure 2 for a visual method overview and Appendix A for further details about experimental setup, implementation, and used prompts.
@@ 2.1 Benchmarks
The NarrativeQA benchmark is a long-document question-answering dataset in English that challenges models to answer questions about stories by reading entire books or movie scripts (Kočiský et al., 2018). We evaluate on the test set, which contains 355 stories (avg. 57K tokens, up to 404K) and 10,557 questions. Each story’s questions are constructed such that high performance requires understanding the underlying narrative, versus relying on shallow pattern matching.
@@ 2.2 Multi-Stage RAG Pipelines
RAPTOR handles long documents by recursively organizing the text into a tree of hierarchical summaries (Sarthi et al., 2024). Concretely, it partitions the text into sentence-level passages, clusters related passages, and uses a language model to summarize each cluster. This process repeats, generating higher-level summaries until a final set of root nodes represents the entire document. At inference time, RAPTOR retrieves from different levels of the summary tree, balancing broad coverage against local detail.
@@ 2.3 Baselines
In our implementation the document is first split into passages capped at 100 tokens, while preserving sentence boundaries where possible. We use neural retrieval with a sentence embedding model (Snowflake Arctic-embed-m 1.5 by Merrick, 2024) to encode both the query and the resulting passages into a shared embedding space. At inference time, passages are ranked by cosine similarity to the query embedding, with the top-ranked passages retrieved until a fixed input token budget (e.g., 10K tokens) is reached. The selected passages, ordered by decreasing similarity, are then concatenated with the query to construct the input to the language model.
@@
Standard RAG pipelines do not preserve the narrative structure within documents, as passages are concatenated solely by retrieval rank. Moreover, retrieval errors can propagate to the downstream language model, which must then reason over potentially missing and disjoint content. To better understand how long-context LMs handle such challenges, we compare against a full-document baseline that simply prompts the model with all available text—eliminating the need to filter passages. We evaluate this baseline only on the QuALITY benchmark, where all documents fit within the language model’s context window.
@@
DOS RAG follows the same retrieval and embedding process as Vanilla RAG, but with one key difference: retrieved passages are reordered to match their original order in the document, not sorted by similarity score. This reordering, achieved by tracking passage positions, preserves original passage order like the full-document baseline while still filtering irrelevant content like Vanilla RAG.
Formally, given a query and a document segmented into passages , let denote the similarity score between and passage . Vanilla RAG retrieves and orders a subset of passages as
In contrast, DOS RAG reorders the same retrieved passages by their original position in the document as
@@ 3 Results
Figure 3 shows performance on the QuALITY benchmark, again with GPT-4o as the reader model. In this setting, we see that all approaches show a steady rise in accuracy as the retrieval budget grows. In particular, full-document baseline with GPT-4o achieves 91.2%, outperforming the best retrieval-augmented systems. Among the retrieval-augmented methods, DOS RAG again achieves the highest performance for token budgets of up to 8K.
@@ 4 Analysis
Why is DOS RAG effective?
We identify four key factors that underlie DOS RAG’s performance and are supported by empirical findings from our evaluation:
1. Retrieving from original passages rather than generated summaries, thereby preserving source information, as in Vanilla RAG and the full-document baseline: We implemented Vanilla RAG as an exact ablation of RAPTOR that excludes generated summaries from the retrieval process. Vanilla RAG consistently outperforms RAPTOR across datasets and retrieval sizes, reinforcing our hypothesis that retrieving directly from original passages results in more robust QA, particularly as long-context LMs reduce the need for intermediate abstraction. While RAPTOR demonstrated superiority in the original paper, its key ablation used UnifiedQA-3B (Khashabi et al., 2020) as the reader model, restricted to a 400-token input. Such a setting highlights RAPTOR’s benefits under constrained context and model capacity, but does not generalize to today’s stronger long-context LMs. Our results with GPT-4o show that, once stronger models and larger context windows are available, retrieving directly from original passages tends to be more robust. This illustrates how advances in model capacity and context length can shift the relative effectiveness of different pipeline designs.
2. Prioritizing retrieval recall over precision while staying within the LM’s effective context size: DOS RAG’s performance increases consistently as the retrieval budget expands up to 30K tokens, after which it plateaus and declines, aligning with prior findings that LMs’ effective context length remains limited (Liu et al., 2024). For shorter documents (6K–8K tokens), the full-document baseline outperforms all methods, indicating that maximizing recall, by including critical information anywhere in the input, can be more effective than precision filtering. However, beyond the effective context window, eliminating irrelevant passages remains essential to maintain performance.
3. Reordering retrieved passages to maintain narrative and argument continuity: Vanilla RAG serves as an exact ablation of DOS RAG, excluding the reordering step. Across all benchmarks and retrieval budgets, DOS RAG consistently outperforms Vanilla RAG, underscoring the benefits of preserving passage order. Performance gain is especially high when the retrieval budget is expanded to tens of thousands of tokens. Retrieving more passages brings us closer to the original document, but without order, the input becomes a disjointed, shuffled version.
4. Favoring simple over complex pipelines: Multi-stage, agentic approaches like ReadAgent decompose QA into multiple LM calls, increasing token usage and latency. However, our evaluation shows that this added complexity does not necessarily improve performance. ReadAgent underperforms compared to DOS RAG at lower token budgets, highlighting the effectiveness of simpler RAG pipelines that use strong embedding models and long-context LMs.
@@ 6 Conclusion
This work examined whether complex multi-stage retrieval pipelines still justify their added complexity given the emergence of long-context LMs capable of processing tens of thousands of tokens. Our controlled evaluation under systematically scaled token budgets shows that simpler methods like DOS RAG can effectively match or even outperform multi-stage pipelines such as ReadAgent and RAPTOR in QA tasks, without intermediate summarization or agentic processing.
We identified four key strategies that contributed to DOS RAG’s performance:
1. Retrieving from original passages rather than generated summaries, maintaining source fidelity and minimizing information loss.
2. Prioritizing retrieval recall over precision, ensuring critical information is included within the effective context window, even at the cost of some less relevant content.
3. Reordering retrieved passages to preserve original passage order, which proved particularly beneficial when dealing with large retrieval budgets.
4. Favoring simple over complex pipelines, while leveraging strong embedding and language models for robustness.
Based on these findings, we recommend establishing DOS RAG as a simple yet strong baseline for future RAG evaluations, paired with state-of-the-art embedding and language models and benchmarked under matched token budgets, to ensure that any added complexity is justified by clear performance improvements as models continue to advance.
@@
Although our results indicate that simpler retrieve-then-read approaches can match or outperform more intricate multi-stage RAG pipelines when paired with long-context language models, our study has several limitations that qualify the generality of these findings.
Our experiments focus on multiple-choice and short-answer reading comprehension tasks over single long documents. We used GPT-4o and GPT-4o-mini as readers and Snowflake’s Arctic-Embed as the embedding model for neural retrieval. While these settings provide useful testbeds for long-context reasoning, it is unclear whether the trends hold for more diverse tasks such as open-ended generation, tasks that require reasoning over multiple documents, or complex reasoning that requires specialized domain knowledge (e.g., in scientific or legal domains). It also remains open whether the findings generalize to other reader LMs (proprietary or open-weight), or alternative retrieval setups with different embedding models. Future work should investigate whether the benefits of simply preserving document continuity extend to these settings, or whether specialized retrieval or summarization steps prove more valuable.
Efficiency is also a key factor in practice. While our comparisons matched token budgets for inference and show DOS RAG competitive across both smaller and larger retrieval windows, we did not measure end-to-end costs of embedding and preprocessing. We estimate that more complex preprocessing, as in RAPTOR and ReadAgent, incurs additional cost, but future work should provide full cost analyses, especially for high-throughput or resource-limited scenarios.
Acknowledgments
We would like to thank the anonymous reviewers for their helpful comments and feedback.
@@ A.7 RAPTOR
We implement RAPTOR using the official repository. To match our other systems, we use the NLTK library for sentence segmentation and the Snowflake Arctic-embed-m 1.5 embedding model (Merrick, 2024) to embed and cluster passages. In all experiments, we use GPT-4o-mini to build the tree of hierarchical summaries to reduce API costs (though note that we experiment with both GPT-4o-mini and GPT-4o as the downstream reader).
@@ Appendix C Full Results
Table 5 presents the results for the NarrativeQA test set across various systems and baselines, using GPT-4o-mini as the reader. Some documents contain up to 404K tokens, far exceeding the 128K context size, which is why we do not report a full-document baseline. Due to issues with the original NarrativeQA download script, three out of 355 stories from the test set were inaccessible, as their document files were empty. Consequently, our results are reported for 352 documents and 10,391 questions for all methods.
@@
| Avg Spent / Budget | F1 | BLEU-1 | BLEU-4 | ROUGE-L | METEOR | |